# Perturbation-based Effect Measures for Compositional Data

### v3

@@ 1.2 Notation
We use superscripts to index coordinates of a vector throughout, that is, for $x\in\mathbb{R}^{d}$ we write $x^{j}$ for the $j$th coordinate of $x$ and denote by $e_{j}$ the $j$th canonical unit vector in $\mathbb{R}^{d}$. For all $n\in\mathbb{N}$, we define $[n]:=\{1,\dots,n\}$. For a real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$, we denote by $\partial_{x^{j}}$ the partial derivative of $f$ with respect to the $x^{j}$ coordinate and the gradient (vector of partial derivatives) of $f$ is denoted by $\nabla f$. When differentiating a function of a single variable at a boundary point of the domain, we only consider the one-sided limit. For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$. Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator.
@@ 2 Target parameters via perturbations
We will discuss how to construct perturbations with interpretable directions and speeds in Section 2.1.2. Choosing to have the direction and speed parametrized by the $1$-norm is an arbitrary choice that turns out to be convenient on the simplex. In this work, we only consider binary and directional perturbations but other classes of perturbations could be of interest too. Perturbations can also be defined on more general sample spaces than simplices and many of the perturbation-related concepts described here can be readily extended to these settings.
@@
Let $\mathcal{Z}\subseteq\Delta^{d-1}$, let $(Y,Z)\in\mathbb{R}\times\mathcal{Z}$ be random variables and define $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$. Then, for a binary perturbation $\psi$, if $f$ is identified on $\{\psi(z^{\prime},1)\in\Delta^{d-1}\mid z^{\prime}\in\mathcal{Z}\}$ and $\mathbb{P}(Z\neq\psi(Z,1))>0$, we define the average (binary) perturbation effect by
@@
Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex (see Section S1 of the supplementary material for a precise definition) at $z$, we define the average (directional) perturbation effect by
@@
We showcase different potential applications for both types of perturbation effect in the numerical experiments in Section 4. For binary perturbations, the denominator is included to scale the effect by the probability that the perturbation does not change $Z$. This is not required but in many cases provides a more meaningful parameter if one is interested in how the perturbation affects points on the simplex that are not yet at their endpoint. For instance, the denominator makes $\lambda_{\psi}$ invariant to modifications of the distribution of $Z$ that only changes $\mathbb{P}(Z\neq\psi(Z,1))$, i.e. the probability that $Z$ is not already perturbed. While both target parameters have obvious similarities to causal effects, they are a priori only (observational) statistical quantities. In particular, they are identifiable from the observational distribution as long as the conditions in Definition 2 are satisfied. We discuss the assumptions required for causal conclusions in Section 2.3.
@@
which is an expression more amenable to analysis by semiparametric theory and thus allows us to construct efficient estimators. If we are willing to assume that $f(\psi(Z,1))-f(Z)$ is constant when $L=0$ and equal to, say $\theta$, then
@@
This derivation is only morally correct when $Z\in\Delta^{d-1}$ ($\nabla f$ is not defined in the usual sense for simplex-valued functions) and to make it formally correct one needs to apply techniques from differential geometry. The proof of the following result can be found in Section S2.1 of the supplementary material.
@@
From this result it follows that if we are able to find a derivative-isolating reparametrization (we describe how to find such reparametrizations for certain perturbations in Section 2.1.2) the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993). It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that
@@
While we do not allow $\omega_{\psi}(z)=0$ in the definition of derivative-isolating reparametrizations, it is clear from the definition of $\tau_{\psi}$ that such points do not contribute to the perturbation effect. We can therefore allow for such points if we appropriately adapt the estimates. In Section S3.2 of the supplementary material, we describe how to modify the estimation procedures introduced in Section 3 to allow for valid estimation even when $\omega_{\psi}(Z)=0$ with positive probability.
@@ 2.1 Perturbations on the simplex
We now turn to the construction of perturbations on the simplex. For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1). For a fixed endpoint there is a canonical (geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint. The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation. We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from an interpretable summary statistic, which decreases when moving away from the endpoint, and interpreting the resulting effect as the effect on the response of slightly increasing the value of the statistic. For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
In the remainder of this section, we discuss interpretable simplex endpoints in Section 2.1.1, provide details on the selection of directions and speed together with the construction of derivative-isolating reparametrizations in Section 2.1.2 and finally in Section 2.1.3 discuss how one could construct directional perturbations that respect a different geometry than the Euclidean one. The perturbations introduced here are used in Section 2.2 to construct specific target parameters based on potential applications. Section 2.2 may be particularly interesting for the reader who is more interested in applications of our framework.
@@ 2.1.1 Endpoints in the simplex
The $(A\to B)$-amalgamation has the effect of setting components in $A$ to 00, scaling up the components in $B$ such that the subcomposition in $B$ remains fixed (i.e., $\mathsf{C}(\mathsf{A}_{A\to B}(z)^{B})=\mathsf{C}(z^{B})$) and leaving the remaining components fixed. We call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto\mathsf{A}_{A\to B}(z)$ amalgamating perturbations. We give an example of a scientific question involving a general amalgamation in Section 2.2.3. An important class of amalgamating perturbations are those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e. perturbations that push towards a vertex of the simplex. Amalgamations, however, do not need to have fixed endpoints but allow for more general endpoints that can also depend on the initial point. Several perturbations on $\Delta^{2}$ are visualized in Figure 2 below.
@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
When constructing directional perturbations it is sufficient to consider perturbations $\psi$ that are parametrized by their direction $v_{\psi}$ and speed $s_{\psi}$ as
$\psi:(z,\gamma)\mapsto z+\gamma s_{\psi}(z)v_{\psi}(z).$
An endpoint function $E:\mathcal{Z}\rightarrow\Delta^{d-1}$ gives rise to a canonical direction defined for all $z\in\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ by
@@
that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint. Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed. In Section 2.1.3 we describe how working in other geometries can lead to different notions of direction. Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect. We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, where the effect of interest corresponds to increasing the value of the statistic, and then constructing a corresponding speed. In both cases, the construction of a derivative-isolating reparametrization is similar.
As a fundamental building block for the general construction below, we consider a perturbation $\psi$ with unit-speed $s_{\psi}=1$ and direction $v_{\psi}=v_{E}$, that is, for all $z\in\Delta^{d-1}$ with $z\neq E(z)$ and all $\gamma\in[0,\|E(z)-z\|_{1}]$ we define
@@
Intuitively, a derivative-isolating reparametrization for $\psi$ should isolate changes along the direction $v_{E}$ from changes due to varying directions. We therefore consider the reparametrization $\phi_{\mathrm{unit}}:\mathcal{Z}\rightarrow\mathbb{R}\times\mathcal{W}$ defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(-\|E(z)-z\|_{1},(E(z),v_{E}(z)))$, where $\mathcal{W}$ is formally defined in Proposition 2. The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=\ell w_{v}+w_{E}$, so
@@
and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$. Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction to start from either a given interpretable speed or a summary statistic. The proof of the following result is given in Section S2.2 of the supplementary material.
Proposition 2 (Finding derivative-isolating reparametrizations).
  1. (a)
    $-s_{w}(\delta)^{-1}=\partial_{\delta}t_{w}(\delta).$ (7)
  2. (b)
    and any other perturbation with the same speed and direction.
Part (a) starts from a given speed and finds a corresponding derivative-isolating reparametrization by solving the differential equation in (7). Part (b) starts from a one-dimensional summary statistic, which is assumed to be smoothly decreasing in the distance from $z$ to the endpoint. Using the statistic as one component in a derivative-isolating reparametrization, the result provides the speed of the perturbation that is being parametrized. We illustrate the use of Proposition 2 by constructing several directional perturbations.
@@
which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition 2 (a). Since for all $w\in\mathcal{W}$ it holds that $w_{E}^{j}=1$ (component $w_{E}$ corresponds to the endpoint $E(z)$) and $w_{v}^{j}=\frac{1}{2}$ (component $w_{v}$ corresponds to the direction $v_{E}(z)$ and we used that $v_{E}(z)^{j}=\frac{1-z^{j}}{\|e_{j}-z^{j}\|_{1}}=\frac{1-z^{j}}{2(1-z^{j})}=\frac{1}{2}$), we have that $\phi_{\mathrm{unit}}^{-1}(\ell,w)^{j}=\frac{\ell}{2}+1$. Therefore,
Hence, solving (7) amounts to solving
$-\frac{2}{(2-\delta)\delta}=\partial_{\delta}t_{w}(\delta)$
@@
Example 4 (Amalgamation influence).
We can extend the considerations of Example 3 to more general amalgamating perturbations. For disjoint $A,B\subseteq[d]$, consider an $(A\to B)$-amalgamating perturbation. Suppose as before that $Z:=\mathsf{C}(X)$ for some $X\in\mathbb{R}^{d}_{+}\setminus\{0\}$ on an absolute scale. We can omit components of $Z$ and $X$ that are not in $A\cup B$ from the analysis as these do not change under the perturbation, so we suppose now that $A\cup B=[d]$. The multiplicative operation described in Example 3 can be generalized by letting $u_{B}=\sum_{j\in B}e_{j}$ and considering
$c\mapsto\mathsf{C}(X\odot(1+cu_{B})),$
$s_{\psi}(z)=\frac{2\|z^{A}\|_{1}\|z^{B}\|_{1}}{(\|z^{A}\|_{1}+\|z^{B}\|_{1})^{2}}.$ (8)
Using Proposition 2 (a) as in Example 3, we further obtain that
@@
Suppose we are interested in the effect of diversifying $z\in\Delta^{d-1}$. We can investigate such an effect by using a centering directional perturbation, that is, a perturbation which satisfies for all $z\in\Delta^{d-1}\setminus\{z_{\mathrm{cen}}\}$ that $v_{\psi}(z)=\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}$. As every element of $\mathcal{W}$ as defined in Proposition 2 is of the form $(z_{\mathrm{cen}},u)$ for a direction $u$, we instead drop the endpoint from the set and use $\mathcal{W}=\{u\in\mathbb{R}^{d}\mid\|u\|_{1}=1,\sum_{j=1}^{d}u^{j}=0\}$ instead. In contrast to the Examples 3 and 4 it is not as obvious how to choose an interpretable speed. We therefore start from a summary statistic and apply Proposition 2 (b) to derive the corresponding speed and derivative-isolating reparametrization. For econometricians a natural choice of diversity measure might be the Gini coefficient. The Gini coefficient $G:\Delta^{d-1}\to[0,1]$ is defined for all $z\in\Delta^{d-1}$ by
$G(z):=\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|z^{j}-z^{k}|.$ (9)
Using $z=z_{\mathrm{cen}}-v_{\psi}(z)\|z_{\mathrm{cen}}-z\|_{1}$, we can express this as
Hence, defining for all $w\in\mathcal{W}$ the strictly decreasing and differentiable function $t_{w}:\mathcal{I}_{w}\rightarrow\mathbb{R}$ ($\mathcal{I}_{w}$ is defined in Proposition 2) for all $\delta\in\mathcal{I}_{w}$ by
$t_{w}(\delta):=-\frac{\delta}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|w^{j}-w^{k}|,$
we have that $t_{v_{\psi}(z)}(\|z_{\mathrm{cen}}-z\|_{1})=-G(z)$. Applying Proposition 2 (b), we conclude that
$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}\right)$
@@
2.1.3 Perturbations arising from alternative simplex geometries
When constructing the directional perturbations in Section 2.1.2, we considered (Euclidean) straight lines to the endpoint to construct directions, that is, $v_{E}:z\mapsto(E(z)-z)/\|E(z)-z\|_{1}$. While we believe that this is natural in many applications, one could also imagine starting from alternative geometries on the simplex and construct other types of directions. We illustrate how this would work for the Aitchison geometry and show that it provides an interesting connection to the log-ratio based literature. The Aitchison geometry is a popular alternative geometry that goes back to the work of Aitchison (1982) who proposed to analyze the interior of the simplex by transforming to a subset of Euclidean space and applying conventional methods to the transformed data. A popular variant of this transformation is the centered log-ratio transform (CLR) defined for all $z\in(\Delta^{d-1})^{\circ}$ (the interior of the simplex) by
or more simply as $\mathrm{CLR}^{-1}(x)=\mathsf{C}(\exp(x))$ where $\exp$ is applied componentwise. The inverse of the $\mathrm{CLR}$ transform can in fact be defined on all of $\mathbb{R}^{d}$ rather than just the range of the $\mathrm{CLR}$-function as is clear from the definition. The $\mathrm{CLR}$-transformation induces what is called the Aitchison geometry (Egozcue and Pawlowsky-Glahn, 2006) on the simplex. The Aitchison geometry induces a different notion of direction than the Euclidean notion considered above and, in particular, the direction of the line connecting $z$ to an endpoint may differ from the direction in (6) (see, for example, Figure 2). However, as $\mathrm{CLR}(z)$ is a bijection on the open simplex, we can just as well consider perturbations on $\mathbb{R}^{d}$ directly and use arguments similar to those in Proposition 2 but now for the perturbation on the log-ratio scale. We illustrate this procedure in the following two examples and discuss connections to the centering and amalgamating perturbations defined above.
Figure 2: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$. The types of perturbation are distinguished by color. The alternative speed for the diversity perturbation in blue is Gini-speed (see (10)) and for the amalgamating perturbations in red and green the speed is multiplicative speed (see (8)). The dashes represent the speed of the perturbation and are equally spaced in the $\gamma$-domain, so that longer lines correspond to faster speeds. The log-ratio diversity perturbation is not a straight line and does not have the same direction as the centering perturbations.
Example 6 (Diversity in the Aitchison geometry).
Example 7 (Feature influence in the Aitchison geometry).

### v4

@@ 1.1 Notation
We use superscripts to index coordinates of a vector throughout, that is, for $x\in\mathbb{R}^{d}$ we write $x^{j}$ for the $j$th coordinate of $x$ and denote by $e_{j}$ the $j$th canonical unit vector in $\mathbb{R}^{d}$. For all $n\in\mathbb{N}$, we define $[n]:=\{1,\dots,n\}$. For a real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$, we denote by $\partial_{x^{j}}$ the partial derivative of $f$ with respect to the $x^{j}$ coordinate and the gradient (vector of partial derivatives) of $f$ is denoted by $\nabla f$. When differentiating a function of a single variable at a boundary point of the domain, we only consider the one-sided limit. For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$. Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator. All references to sections and results beginning with ‘S’ point to the supplementary material.
@@ 2 Target parameters via perturbations
We construct interpretable perturbations in Section 2.1.2. Choosing to have the direction and speed parametrized by the $1$-norm is an arbitrary choice that turns out to be convenient on the simplex. The concept of a perturbation can be defined in greater generality than we do here but we stick to the simplex for simplicity.
@@
Let $\mathcal{Z}\subseteq\Delta^{d-1}$, let $(Y,Z)\in\mathbb{R}\times\mathcal{Z}$ be random variables and define $f:z\mapsto\mathbb{E}[Y\,|\,Z=z]$. Then, for a binary perturbation $\psi$, if $\mathbb{P}(Z\neq\psi(Z,1))>0$ and $z\mapsto\psi(z,1)$ preserves null sets of $Z$, we define the average (binary) perturbation effect by
@@
Furthermore, for a directional perturbation $\psi$, if for all $z\in\mathcal{Z}$ it holds that $f$ is differentiable on the simplex (see Section S1 for a precise definition) at $z$, we define the average (directional) perturbation effect by
@@
The conditions we impose on the perturbation and distribution of $Z$ are only to ensure that the parameters are well-defined (see Section S2.3). We showcase different potential applications for both types of perturbation effect in the examples in Section S2.2 and the numerical experiments in Section 4. For binary perturbations, the denominator is included to scale the effect by the probability that the perturbation does not change $Z$. This is not required but in many cases provides a more meaningful parameter if one is interested in how the perturbation affects points on the simplex that are not yet at their endpoint. For instance, the denominator makes $\lambda_{\psi}$ invariant to modifications of the distribution of $Z$ that only changes $\mathbb{P}(Z\neq\psi(Z,1))$, i.e. the probability that $Z$ is not already perturbed. While both target parameters have obvious similarities to causal effects, they are a priori only (observational) statistical quantities. We discuss the assumptions required for causal conclusions in Section S3.
@@
which is an expression more amenable to analysis by semiparametric theory and thus allows us to construct efficient estimators. If we are willing to assume that $f(\psi(Z,1))-f(Z)$ is constant when $L=0$ and equal to $\theta$, then
@@
This derivation is only morally correct when $Z\in\Delta^{d-1}$ ($\nabla f$ is not defined in the usual sense for simplex-valued functions) and to make it formally correct one needs to apply techniques from differential geometry. The proof of the following result can be found in Section S2.4.
@@
From this result it follows that if we are able to find a derivative-isolating reparametrization the average directional perturbation effect can be expressed as an average partial derivative of a conditional expectation (4) – a quantity that has been previously studied using semiparametric theory (Newey and Stoker, 1993). It can either be estimated nonparametrically or by assuming a partially linear model, that is, assuming that there exists a parameter $\theta\in\mathbb{R}$ and a function $h$ such that
@@
While we do not allow $\omega_{\psi}(z)=0$ in the definition of derivative-isolating reparametrizations, it is clear from the definition of $\tau_{\psi}$ that such points do not contribute to the perturbation effect. We can therefore allow for such points if we appropriately adapt the estimates. In Section S4.4, we describe how to modify the estimation procedures introduced in Section 3 to allow for valid estimation even when $\omega_{\psi}(Z)=0$ with positive probability.
@@ 2.1 Perturbations on the simplex
We now turn to the construction of perturbations on the simplex. For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1). For a fixed endpoint there is a canonical (geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint. The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation. We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from an interpretable summary statistic, which decreases when moving away from the endpoint, and interpreting the resulting effect as the effect on the response of slightly increasing the value of the statistic. For both approaches, we provide – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
@@ 2.1.1 Endpoints in the simplex
The $(A\to B)$-amalgamation has the effect of setting components in $A$ to 00, scaling up the components in $B$ such that the subcomposition in $B$ remains fixed (i.e., $\mathsf{C}(\mathsf{A}_{A\to B}(z)^{B})=\mathsf{C}(z^{B})$) and leaving the remaining components fixed. We call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto\mathsf{A}_{A\to B}(z)$ amalgamating perturbations. We discuss general amalgamating perturbations in Section S2.6. Here, we focus on an important class of amalgamating perturbations: those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e. perturbations that push towards a vertex of the simplex.
@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
When constructing directional perturbations it is sufficient to consider perturbations $\psi$ that are parametrized by their direction $v_{\psi}$ and speed $s_{\psi}$ as $\psi:(z,\gamma)\mapsto z+\gamma s_{\psi}(z)v_{\psi}(z).$ An endpoint function $E:\mathcal{Z}\rightarrow\Delta^{d-1}$ gives rise to a canonical direction defined for all $z\in\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ by
@@
that is, for each $z$ we define the unit vector $v_{E}(z)$ in the direction of the straight line from $z$ to the endpoint. Throughout this section, to avoid zero speeds, we restrict the domain of the perturbations to $\{z\in\Delta^{d-1}\mid E(z)\neq z\}$ but as discussed in Remark 1 the target parameters and estimates can easily be extended to settings where zero speeds are allowed. In Section S2.7 we describe how working in other geometries can lead to different notions of direction. Taking the direction in (6) as given, we now want to select a speed $s_{\psi}$ that leads to an interpretable average directional perturbation effect and then construct a derivative-isolating reparametrization for the resulting perturbation in order to efficiently estimate the effect. We propose two approaches: (a) starting directly from a speed that satisfies some regularity conditions and (b) starting from a one-dimensional summary statistic, where the effect of interest corresponds to increasing the value of the statistic, and then constructing a corresponding speed. In both cases, the construction of a derivative-isolating reparametrization is similar.
As a fundamental building block for the general construction, we consider a perturbation $\psi$ with unit-speed $s_{\psi}=1$ and direction $v_{\psi}=v_{E}$, that is, for all $z\in\Delta^{d-1}$ with $z\neq E(z)$ and all $\gamma\in[0,\|E(z)-z\|_{1}]$ we define
@@
Intuitively, a derivative-isolating reparametrization for $\psi$ should isolate changes along the direction $v_{E}$ from changes due to varying directions. We therefore consider the reparametrization defined for all $z\in\mathcal{Z}$ by $\phi_{\mathrm{unit}}(z):=(-\|E(z)-z\|_{1},(E(z),v_{E}(z)))$. The inverse of this reparametrization is $\phi_{\mathrm{unit}}^{-1}(\ell,(w_{E},w_{v}))=\ell w_{v}+w_{E}$, so
@@
and hence $\phi_{\mathrm{unit}}$ is derivative-isolating for $\psi$. Since a unit-speed perturbation may not always be easy to interpret, we extend this basic construction in Section S2.5 to start from either a given interpretable speed or a summary statistic. We illustrate both options in the following two examples.
@@
which we call multiplicative speed. To find a corresponding derivative-isolating reparametrization we apply Proposition S1 (a) (see Example S3 for details) and obtain that
@@
Suppose we are interested in the effect of diversifying $z\in\Delta^{d-1}$. We can investigate such an effect by using a centering directional perturbation, that is, a perturbation which satisfies for all $z\in\Delta^{d-1}\setminus\{z_{\mathrm{cen}}\}$ that $v_{\psi}(z)=\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}$. In contrast to Example 1 it is not as obvious how to choose an interpretable speed. For econometricians a natural choice of similarity measure might be the Gini coefficient. The Gini coefficient $G:\Delta^{d-1}\to[0,1]$ is defined for all $z\in\Delta^{d-1}$ by
$G(z):=\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|z^{j}-z^{k}|,$ (7)
whence we define a diversity measure by $D(z):=1-G(z)$. Using Proposition S1 (b) (see Example S4 for details), we obtain that
$\phi=(\phi^{L},\phi^{W}):z\mapsto\left(1-G(z),\frac{z_{\mathrm{cen}}-z}{\|z_{\mathrm{cen}}-z\|_{1}}\right)$
@@
Figure S1 contains an illustration of the different perturbations we have considered in this work.